# Symbolic Control for Autonomous Docking of Marine Surface Vessels

### v1

@@ IV Experimental Setup
We test the correct-by-construction hierarchical control architecture in performing autonomous docking maneuvers through simulation and real-world experiments.
IV-A Simulation
@@
where $\Delta t$ represents the time step. Hydrodynamic coefficients for $\mathbf{M}$, $\mathbf{C}(\nu)$, and $\mathbf{D}(\nu)$ were obtained via system identification in [30]. The simulation environment is implemented using Robot Operating System 2 (ROS 2) and runs in real-time. It models the vessel’s surge, sway, and yaw motions by integrating the forces and torques applied by multiple thrusters, utilizing the integrator in (7). The simulator updates the vessel’s position ($\eta$) and velocity ($\nu$) based on thrust commands received from ROS topics. It then publishes motion data in common ROS interfaces such as odometry and pose.
IV-B Towing Tank Environment
@@
1m1m1m3m50cm8m6m
(b) Test environment size
(c) Cybership Voyager
Figure 7: LABEL:sub@fig:arena_picture Marine Cybernetics Laboratory where experiments took place in a towing tank ($40m\times 6.45m\times 1.5m$) with two yellow, two stationary obstacles ($0.5m\times 3m$) attached to the walls and a scaled model vessel (length $1m$). LABEL:sub@fig:arena_specs Environment dimensions LABEL:sub@fig:voyager Model vessel that was used in the experiments
Experimental trials were conducted in NTNU’s Marine Cybernetics Laboratory (MCLab) using a scaled model vessel, the C/S Voyager, in its towing tank. The towing tank has dimensions of $40m\times 6.45m\times 1.5m$, and is equipped with a real-time positioning system, Qualisys, which includes Oqus cameras and the Qualisys Track Manager software. This environment is ideal for testing motion control systems for marine vessels, due to its manageable size and advanced instrumentation.
The main test involved two stationary obstacles for the vessel to evade while approaching the target docking position. The testing region for this experiment was a $8m\times 6m\times 1.5m$ portion of the towing tank. In this region, we used 7 motion capturing cameras, ensuring full coverage of the testing area. The two stationary obstacles, each $0.5m\times 3m$, were attached to the walls of the towing tank as depicted in Fig. 7a and 7b.
IV-C Model Vessel
The C/S Voyager, shown in Fig. 7c, is a 1:32 scale tugboat built at NTNU’s MCLab to test advanced navigation and control systems. It is highly maneuverable due to its fast-responding thrusters. The C/S Voyager’s azimuth thrusters can reach a maximum force of more than 80N, and it has over 4 hours of endurance. Real-time communication between the vessel’s components is made possible by ROS 2, which is installed on a Raspberry Pi 4, the onboard computer. The C/S Voyager has an array of sensing devices, including infrared spheres for accurate position tracking using the Qualisys motion capture system and Inertial Measurement Units (IMUs) for measuring linear accelerations and angular velocity rates. Dimensions of the C/S Voyager are given in Fig. 8.
[…]
Figure 8: C/S Voyager dimensions, $l_{y1}=l_{y2}=7.0cm$, $l_{x1}=l_{x2}=41.5cm$, $l_{x3}=37.0cm$, $l_{LOA}=100.0cm$, and $l_{b}=33.0cm$.
IV-D Optimization Tuning
Since the controller synthesis utilizes a simplified mathematical model of the vessel dynamics, some of the synthesized actions do not correspond to reasonable vessel behavior. For example, it may be safe for the vessel to traverse the docking path backwards; however, it is impractical to exhibit this behavior in the physical world. The matrix $W$, introduced in Section III-A, ensures that only reasonable actions are chosen.
We designed matrix $W$ to ensure the optimal control action is chosen from the synthesized action set. To determine the value of $W$, we performed empirical tests and choose the $W$ that provided the smoothest and most realistic vessel behavior. $W$ is chosen conditionally based on the forward velocity component of the proposed control action, $\sigma_{i_{\nu_{x}}}$, such that
@@ IV-D Optimization Tuning
Given a proposed action with a negative forward velocity, we heavily penalize this backward motion, with weight 7.5, discourage sway motion, with weight 3, and minimize drastic changes in the yaw, with weight 2.5. This ensures that the vessel prioritizes forward motion and limits the usage of the tunnel thruster. Additionally, (8) minimizes rapid rotations of the vessel by maintaining the heading of the vessel, making sure it does not drastically change in a single time step. When a proposed action has a non-negative forward velocity, we only penalize sway motion and changes in the yaw. However, we must increase the sway penalty, with weight 6, as the synthesized controller is more likely to suggest sway commands when we are making forward progress. We similarly discourage vessel rotations and changes in the yaw, with weight 2.5.
@@ V Results
We executed 2 types of tests on the simulator: experimental validation and dynamic obstacles. First, we mirrored the experimental setup, shown in Fig. 7a, to test the efficacy of our control strategy on both a single obstacle and two obstacles before deploying it on the real vessel. Because we were able to successfully synthesize a control strategy that safely traversed obstacles in simulation, we quickly begin real-world testing.
Additionally, we tested the adaptability of symbolic control in the presence of a dynamic obstacle in the simulation environment. Without adding extra layers of control, we successfully conducted 10 autonomous docking maneuvers while evading a dynamical obstacle. An example of the vessel’s behavior in this scenario is shown in Fig. 9. In these tests, the obstacle, sized at $1m\times 1m$, started from the position $(X:0.5,Y:-2.5)$ and moved at a constant velocity of $0.15m/s$ in the positive $Y$ direction. The vessel followed its normal trajectory until the obstacle entered its vicinity at $T=15s$, at which point the vessel stopped moving until the obstacle had cleared the way before resuming its approach to the docking area, denoted by “End” [^1] Video of simulated docking maneuvers: https://www.youtube.com/watch?v=6eLPi4lUYQ8.
Since we did not control the obstacle’s path, safety was only ensured against indirect collisions. When the obstacle moved in a path that could lead to a side or head-on collision, avoiding impact was sometimes impossible. This limitation was confirmed through simulations where the obstacle’s trajectory directly opposed the vessel’s intended docking maneuver. We leave further testing, and refinement of the dynamic obstacle’s behavior, as future work.
Figure 9: Example trajectory of a successful autonomous docking in which the vessel evades a square dynamic obstacle. The trajectory of the vessel is shown at various time steps ($t=5,15,25,35$), along with the position of the dynamic obstacle at that time, to show how the vessel holds a safe position until the obstacle has left the vicinity.
All implementations for the experimental validation were implemented using ROS 2. The necessary subsystems, such as the observers and DP controllers described in Fig. 6, were run on the onboard single board computer. The correct-by-construction synthesis task, as detailed in Section III-A, was performed on a server every two seconds given a state set reflecting this setup as shown in Fig. 1. The resulting velocity commands are guaranteed to maintain vessel safety, given the provable nature of symbolic control.
@@
Figure 10: Trajectories (gray) of 10 successful autonomous docking maneuvers. There was a range of vessel starting positions, the yellow rectangles portray the 3 meter obstacles, and the green box denotes the 1 meter square target.
@@
The trajectories of our successful autonomous docking maneuvers are shown in Fig. 10. The vessel maneuvered generously around the obstacles, allowing extra space to maintain safety and ensure there would be no collisions [^2] Video of successful autonomous docking maneuver in MCLab with model vessel: https://www.youtube.com/watch?v=q-qohEJciU4. On average, the vessel completed this trajectory in 3.41 minutes.
@@ VI Discussion and Future Work
We executed the real-world two-obstacle experiment 25 times, resulting in 10 successful autonomous docking maneuvers. We recognize four categories of experimental issues: communication errors, localization errors, optimization tuning challenges, and synthesis adjustments. During the real-world experiments, network errors led to communication drops lasting between 5 and 20 seconds. Because position measurements were transmitted to the vessel over the network, these interruptions significantly impacted observer accuracy. The synthesis server continued to command new velocities without updated position data, leading to either incomplete runs or erratic vessel behavior due to missing velocity commands. In addition to the network errors, the server was crashing without any error message due to constant polling by the vessel to catch the synthesis state changes. This issue caused experimental failure, discovered only after completion of testing.
Further, we analyzed the ability of the low-level controller to realize the synthesized velocity commands during each successful experiment. To quantify the error between the commanded and realized velocities, we calculated the mean squared error (MSE) of $\nu_{\text{\tiny{MSE}}}=[0.00768\nicefrac{{m}}{{s}},0.00431\nicefrac{{m}}{{s}},0.62108\nicefrac{{deg}}{{s}}]^{\top}$. How this error in the low-level controller affects the formal guarantees of the proposed approach requires future investigation. We note that this error analysis hints to the robustness of our approach. Even when correctness assumptions on the low-level controller were not met, the accuracy of the low-level controller did not affect the efficacy of the upper-level symbolic control, in many cases.
This system was created with several limitations: it utilizes a simplified vessel model, and it neglects environmental loads, such as wind and waves. In future work, we hope to address these limitations by developing the current vessel model through physics-informed system identification and introducing wave and wind models. Additionally, while we demonstrated the ability of this method to adapt to specific dynamic obstacles without any modifications in simulation, we recognize the need to analyze and test this further in real-world situations. Finally, there are additional applications of this hierarchical symbolic control scheme that could be explored. For example, we will consider an optimization methodology through the development of an energy management system to provide a correct-by-construction control architecture that solves the optimal load sharing problem for vessels.
@@ VII Conclusion
In this paper we have developed a hierarchical symbolic control strategy for the safe autonomous docking of a DP vessel. To validate and verify the proposed design, simulation and experimental trials were presented. The simulation shows the ability of this approach to adapt to new environmental conditions, such as dynamic obstacles, without requiring further levels of control to be developed. Furthermore, the real-world experimental results demonstrate that the vessel can safely traverse a docking path given stationary obstacles.
@@ Acknowledgments
This paper is supported in part by the NSF Cyber-Physical Systems Frontier project CNS-2111688, NTNU VISTA Centre for Autonomous Robotic Operations Subsea (CAROS), Research Council of Norway (RCN) through SFI AutoShip (RCN project 309230), Guangzhou-HKUST(GZ) Joint Funding Program (Grant No.2023A03J0008), Education Bureau of Guangzhou Municipality, and a grant from the Peder Sather Center for Advanced Study at UC Berkeley. The first author was also supported by a NSF Graduate Research Fellowship. The authors thank Emil Bratlie, Robert Opland, Terje Rosten, Vebjørn Steinsholt, Gabriel Bjørgan Vikan, and Anne Elise Havmo for facilitating experimentation in the MCLab.

### v2

@@ IV Experimental Setup
IV-A Problem Formulation
We refine the definition of our symbolic control problem (3) based on preliminary experimental observations. With the goal of ensuring reliable performance, we tracked the response time of our low-level velocity controller and defined the set of possible control inputs to be the best-performing combinations of linear and angular velocities while accounting for the testing arena and vessel capabilities. This resulted in the set of inputs, $U_{1}$, and vessel states, $X_{1}$, defined as,
$\begin{aligned} U_{1}&=[-0.1,0.2]\times[-0.1,0.1]\times[-0.2,0.2]\\ X_{1}&=[-3.5,4.5]\times[-3.0,3.0]\times[-\pi,\pi).\end{aligned}$ (5)
IV-B Simulation
@@
where $\Delta t$ represents the time step. Hydrodynamic coefficients for $\mathbf{M}$, $\mathbf{C}(\nu)$, and $\mathbf{D}(\nu)$ were obtained via system identification in [32]. The simulation environment is implemented using Robot Operating System 2 (ROS 2) and runs in real-time. It models the vessel’s surge, sway, and yaw motions by integrating the forces and torques applied by multiple thrusters, utilizing the integrator in (6). The simulator updates the vessel’s position ($\eta$) and velocity ($\nu$) based on thrust commands received from ROS topics and publishes motion data in common ROS interfaces.
IV-C Towing Tank Environment
Experimental trials were conducted in NTNU’s Marine Cybernetics Laboratory (MCLab) using a scaled model vessel, the C/S Voyager. The towing tank has dimensions of $40m\times 6.45m\times 1.5m$, and is equipped with a real-time positioning system, Qualisys, which includes Oqus cameras and the Qualisys Track Manager software. This environment is ideal for testing motion control systems for marine vessels, due to its manageable size and advanced instrumentation.
The main test involved two stationary obstacles for the vessel to evade while approaching the target docking position. The testing region for this experiment was a $8m\times 6m\times 1.5m$ portion of the towing tank. In this region, we used 7 motion capturing cameras, ensuring full coverage of the testing area. The two stationary obstacles, each $0.5m\times 3m$, were attached to the walls of the towing tank as depicted in Fig. 6a. No external forces (e.g. waves) were introduced to the tank during testing.
@@ IV-C Towing Tank Environment
(b) C/S Voyager
Figure 6: LABEL:sub@fig:arena_picture Marine Cybernetics Laboratory (MCLab) where experiments took place in a towing tank ($40m\times 6.45m\times 1.5m$) with two yellow, stationary obstacles ($0.5m\times 3m$) attached to the walls and a scaled model vessel (length $1m$).LABEL:sub@fig:voyager Model vessel that was used in the experiments
IV-D Model Vessel
The C/S Voyager, shown in Fig. 6b, is a 1:32 scale tugboat built at NTNU’s MCLab to test advanced navigation and control systems. The C/S Voyager has over 4 hours of endurance and azimuth thrusters, which can reach a maximum force of more than 80N. The vessel’s components achieve real-time communication through ROS 2, which is installed on the onboard computer, a Raspberry Pi 4. The C/S Voyager has an array of sensing devices, including infrared spheres for accurate position tracking with the Qualisys motion capture system and Inertial Measurement Units (IMUs) for measuring linear accelerations and angular velocities.
IV-E Optimization Tuning
Since the controller synthesis utilizes a simplified vessel dynamics model, some of the synthesized actions may not reflect realistic vessel behavior. For example, it may be safe for the vessel to perform docking backwards; however, it is impractical to exhibit this behavior in the physical world. The matrix $W$, introduced in Section III-A, ensures that only reasonable actions are chosen. We designed $W$ to ensure the optimal control action was chosen from the synthesized action set. We derive $W$ systematically using domain specific knowledge and design iterations. $W$ is chosen conditionally based on the forward velocity component of the proposed control action, $\sigma_{i_{\nu_{x}}}$, such that
@@ IV-E Optimization Tuning
Given a proposed action with a negative forward velocity, we penalize this backward motion (weight 7.5), discourage sway motion (weight 3), and minimize drastic changes in the yaw (weight 2.5). This ensures that the vessel prioritizes forward motion and limits usage of the tunnel thruster. Additionally, (7) minimizes rapid rotations of the vessel, reducing drastic changes in the heading. When a proposed action has a non-negative forward velocity, we only penalize sway motion and changes in the yaw. However, we must increase the sway penalty (weight 6), as the synthesized controller is more likely to suggest sway commands when we make forward progress. We similarly discourage vessel rotations and changes in the yaw (weight 2.5).
@@ V Results
To test the efficacy of our control strategy on both one and two obstacles before deploying to the real vessel, we performed experimental trials in simulation, mirroring the setup shown in Fig. 6a. We were able to successfully synthesize control strategies that safely traversed static obstacles in simulation, as seen in Fig. 7a  [^1] Docking in simulation environment: https://youtu.be/TFynNTElZV8.
@@
Figure 7: The yellow rectangles portray 3 meter obstacles, and the green box denotes the 1 meter square target. LABEL:sub@fig:sim_run Trajectories of 5 successful docking maneuvers in the simulation environment, all starting from the same vessel position. LABEL:sub@fig:successful_runs Trajectories of 10 successful autonomous docking maneuvers in the real-world, starting from a range of vessel positions.
To execute symbolic controller synthesis in real-time, we utilized the pFaces [21] accelerator deployed on a server equipped with an Nvidia A6000 GPU (48 GB of VRAM), 256 GB of memory, and an Intel Xeon 6226R processor with the pFaces Symbolic Control kernel [33]. The server performs the correct-by-construction synthesis task, as detailed in Section III-A, in approximately 0.5 seconds, enabling us to perform controller re-synthesis in real-time. To enable real-time interaction with the autonomous vessel, we developed a ROS 2 Python package that served as an interface to communicate control commands and vessel state. This architecture, as outlined in Fig. 1, provides a robust and scalable computational infrastructure for the iterative development, validation, and real-time implementation of advanced control strategies.
State information was relayed to the server after every completed synthesis task. The server allocated apx. 400MB of VRAM and apx. 500MB of system memory to complete the synthesis task. The necessary subsystems, such as the observers and DP controllers described in Fig. 5, were run on the onboard single board computer, a Raspberry Pi 4 with 8GB memory. The symbolic controller assumes: (1) the low-level control system is able to realize any commanded velocity and (2) the kinematic model, which the symbolic control is based upon, encompasses all possible behaviors of the real-world vessel. We observed this to be true in simulation, Eq. (6), and during the real-world experiments. Given these assumptions, the synthesized velocity commands are theoretically guaranteed to maintain vessel safety. We discuss discrepancies between these theoretical guarantees and the practical implementation in Section VI.
@@
The trajectories of the successful autonomous docking maneuvers, completed in the laboratory environment with the scale model vessel, are shown in Fig. 7b222 Docking of model vessel in towing tank: https://youtu.be/q-qohEJciU4. The vessel maneuvered generously around the obstacles, allowing extra space to maintain safety and ensure there would be no collisions. On average, the vessel completed this trajectory in 2.67 minutes; the minium and maximum durations were 1.71 and 3.66 minutes, respectively.
@@ VI Discussion and Future Work
We analyzed the low-level controller’s ability to realize the synthesized velocity commands. To quantify the error between commanded and realized velocities, we calculated the mean squared error (MSE), $\nu_{\text{\tiny{MSE}}}=[0.00768\nicefrac{{m}}{{s}},0.00431\nicefrac{{m}}{{s}},0.62108\nicefrac{{deg}}{{s}}]^{\top}$. Although we did not formally account for the tracking error, it is possible to do so by performing a continuous abstraction in which the low-level controller guarantees a bound on the tracking error and the high-level symbolic synthesis accounts for the error, as proposed in [23]. Replacing the existing low-level DP controller in experiments with one obtained from a continuous abstraction is left for future research. Additionally, we employed a simplified vessel model, which neglects environmental loads, such as wind and waves, and we relied on vessel-server communication. In future work, we hope to address these limitations by developing the current vessel model through physics-informed system identification, introducing wave and wind models, and advancing the hardware capabilities of the C/S Voyager to enable experimental testing in outdoor environments. Further, we hope to investigate the extension of this approach to full-scale vessels.
@@ VII Conclusion
We presented a hierarchical symbolic control strategy for safe autonomous docking. We utilized parallel computing and GPU acceleration to achieve real-time synthesis in dynamic settings, with experimental validation marking a key step toward real-world implementation. The work represents a new level of maturity in symbolic control, demonstrating real-time synthesis despite past computational limitations.
@@ Acknowledgments
This paper is supported in part by the NSF project CNS-2111688, NTNU VISTA Centre for Autonomous Robotic Operations Subsea (CAROS), Research Council of Norway (RCN) through SFI AutoShip (RCN project 309230), Guangzhou-HKUST(GZ) Joint Funding Program (Grant No.2025A03J4493), Education Bureau of Guangzhou Municipality, and a grant from the Peder Sather Center for Advanced Study at UC Berkeley. The first author was also supported by a NSF Graduate Research Fellowship. The authors thank Emil Bratlie, Robert Opland, Terje Rosten, Vebjørn Steinsholt, Gabriel Bjørgan Vikan, and Anne Elise Havmo of the MCLab.